OpenStudy (awkwardpanda):

if cos theta = 0.5, find sec theta. 0.25 <--- 0.5 1 2

OpenStudy (bibby):

what is secant?

OpenStudy (awkwardpanda):

the ratio of the hypotenuse to the shorter side adjacent to an acute angle. right?

OpenStudy (bibby):

kinda. in that cos = adj/hyp sec = hyp/adj secant is 1/cos

OpenStudy (bibby):

so if \(cos(\theta)=\frac{5}{3}\) \(\huge sec(\theta)=\frac{1}{\frac{5}{3}}=\frac{3}{5}\)

OpenStudy (awkwardpanda):

3/5 comes out to 0.6, but 0.6 isn't in my choices.

OpenStudy (awkwardpanda):

my head hurts

OpenStudy (bibby):

I made up an example to demonstrate the concept

OpenStudy (bibby):

if cos = .5 sec = 1/0.5=\(\large \frac{1}{\frac{1}{2}}=\frac{1}{}*\frac{2}{1}=2\)

OpenStudy (anonymous):

cos (theta)=1/sec(theta)

OpenStudy (anonymous):

1/2=1/(1/2)=2

OpenStudy (awkwardpanda):

you all are amazing. im cramming like 2 projects rn and im gonna die. thank you for the help

OpenStudy (bibby):

I hope your head stops sucking. Good luck dude

OpenStudy (awkwardpanda):

me too, thank you.